OpenStudy (anonymous):

7(9x - 3) - 4(2x - 1) =

OpenStudy (anonymous):

7(9x - 3) = 63x - 21 -4(2x - 1) = -8x + 4 (63x - 21) - (-8x + 4) = 71x - 25

OpenStudy (anonymous):

alright i think im getting this.. :p thanks ;)